OpenStudy (anonymous):

what is the slope of a line parallel to the line 5x-9y=1?

OpenStudy (anonymous):

5x-9y=1 -9y=1-5x -9y=1-5x y=-1/9+(5/9)x slope of the line is 5/9 so the line parallel to this line must also have a slope of 5/9